OpenStudy (michaelpersaus48):

HELP!!! What is the equation of the perpendicular bisector of the segment joining points A(2, 2) and B(6, 8)?

OpenStudy (will.h):

1st we need to find the midpoint of the coordinates above (2,2)(6,8) midpoint : x1 + x2/2 , y1+y2/2) midpoint: (4,5) now we need to find the slope of the perpendicular line which is the slope of the reciprocal of the original coordinates Slope = 6/4 Slope = 3/2 the reciprocal for that is -2/3 Now we have the slope as -2/3 and the line would pass through (4,5) use y - y1 = m(x - x1) m is the slope substitute to get the equation of the perpendicular line y - 5 = -2/3(x - 4) y - 5 = (-2/3)x + 8/3 y = (-2/3)x + 23/3

OpenStudy (michaelpersaus48):

thank you bro!!!

OpenStudy (will.h):

No problem